# Powerformer: A Transformer with Weighted Causal Attention for Time-series Forecasting

### v1

@@ Acknowledgements
Figure A23: We present the Powerformer attention score (inset) and weight distributions on the ETTh1 dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A24: We present the Powerformer attention score (inset) and weight distributions on the ETTh1 dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A25: We present the Powerformer attention score (inset) and weight distributions on the ETTh2 dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A26: We present the Powerformer attention score (inset) and weight distributions on the ETTh2 dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A27: We present the Powerformer attention score (inset) and weight distributions on the ETTm1 dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A28: We present the Powerformer attention score (inset) and weight distributions on the ETTm1 dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A29: We present the Powerformer attention score (inset) and weight distributions on the ETTm2 dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A30: We present the Powerformer attention score (inset) and weight distributions on the ETTm2 dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
@@
Figure A31: We present the Powerformer attention score (inset) and weight distributions on the Weather dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
@@
Figure A32: We present the Powerformer attention score (inset) and weight distributions on the Weather dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
@@
Figure A33: We present the Powerformer attention score (inset) and weight distributions on the Electricity dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
@@
Figure A34: We present the Powerformer attention score (inset) and weight distributions on the Electricity dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A35: We present the Powerformer attention score (inset) and weight distributions on the Traffic dataset with a forecast and input length of 96 and 336, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A36: We present the Powerformer attention score (inset) and weight distributions on the Traffic dataset with a forecast and input length of 720 and 336, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
A6.2.2 Similarity Power-Law Mask
@@
Table A6 presents the aggregate MSE and MAE results for the similarity power-law mask ($f^{(\text{SPL})}(t)$) with varying time decays ($\alpha$), forecast lengths, and input sequence lengths. The 512 input sequence generally outperforms the shorter 336 input length. Compared to $f^{(\text{PL})}(t)$, there are more instances in which a 336 input sequence length outperforms. This is likely due to the much faster decay of $f^{(\text{SPL})}(t)$ compared to $f^{(\text{PL})}(t)$. As expected, the best performing $\alpha$ varies between datasets. However, there is much less variation as a function of forecast length and input sequence length when compared to $f^{(\text{PL})}(t)$.
@@
We compare all datasets’ attention score and weight distributions for weighted causal attention in Figs A37-A50. Due to the faster decay of $f^{(\text{SPL})}(t)$ we observe stronger bimodal distributions for all datasets than for $f^{(\text{PL})}(t)$, with wider $\mathbf{C}^{(\text{C,D})}$ distributions for ETTm1, ETTm2, and Weather.
@@
Table A6: We compare Powerformer’s performance with the similarity power-law mask $f^{(\text{SPL})}(t)$ on standard time-series datasets for varying decay lengths. The best results are bolded and the second best are underlined.
@@
Figure A37: We present Powerformer’s attention score and weight distributions on the ETTh1 dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A38: We present Powerformer’s attention score and weight distributions on the ETTh1 dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A39: We present Powerformer’s attention score and weight distributions on the ETTh2 dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A40: We present Powerformer’s attention score and weight distributions on the ETTh2 dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A41: We present Powerformer’s attention score and weight distributions on the ETTm1 dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A42: We present Powerformer’s attention score and weight distributions on the ETTm1 dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A43: We present Powerformer’s attention score and weight distributions on the ETTm2 dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A44: We present Powerformer’s attention score and weight distributions on the ETTm2 dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
@@
Figure A45: We present Powerformer’s attention score and weight distributions on the Weather dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
@@
Figure A46: We present Powerformer’s attention score and weight distributions on the Weather dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
@@
Figure A47: We present Powerformer’s attention score and weight distributions on the Electricity dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
@@
Figure A48: We present Powerformer’s attention score and weight distributions on the Electricity dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A49: We present Powerformer’s attention score and weight distributions on the Traffic dataset with a forecast and input length of 96 and 336, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
Figure A50: We present Powerformer’s attention score and weight distributions on the Traffic dataset with a forecast and input length of 720 and 336, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents WCMHA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,D})}$, and the solid lines represent WCMHA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,D})}$.
A6.2.3 Butterworth Filter: Order 1
@@
Here we evaluate the Butterworth order 1 mask ($f^{(\text{BW})}_{1}(t)$) on the ETT and Weather datasets with a 336 look-back window. The $f^{(\text{BW})}_{1}(t)$ decay is not as steep as $f^{(\text{PL})}(t)$ but has a more gradual decay than $f^{(\text{BW})}_{2}(t)$, as shown in Figs. A1 and 2. This mask offers a middle ground between the step function like $f^{(\text{BW})}_{2}(t)$ and the faster decaying $f^{(\text{PL})}(t)$. Table A7 presents the results for Powerformer with WCMHA and $f^{(\text{BW})}_{1}(t)$. We observe that $f^{(\text{BW})}_{1}(t)$ underperforms $f^{(\text{PL})}(t)$, $f^{(\text{SPL})}(t)$ and $f^{(\text{BW})}_{2}(t)$. The poor performance further supports our hypothesis that the functional form of $f(t)$ is important when imposing a locality bias.
@@
Table A7: We compare Powerformer’s performance with WCMHA and the order 1 Butterworth Filter mask $f^{(\text{BW})}_{1}(t)$ on standard time-series datasets for varying decay lengths. The best results are bolded and second best are underlined.
@@
A6.2.4 Butterworth Filter: Order 2
Here we evaluate the Butterworth order 2 mask ($f^{(\text{BW})}_{2}(t)$) on the ETT and Weather datasets with a 336 look-back window. The $f^{(\text{BW})}_{2}(t)$ decay is analogous to the step function, but has a smooth decay to 0 at the end of the window, as shown in Figs. A1. We use $f^{(\text{BW})}_{2}(t)$ as an improved step function to avoid the sharp cutoff. Table A8 presents the results for Powerformer with WCMHA and $f^{(\text{BW})}_{2}(t)$. We observe that $f^{(\text{BW})}_{2}(t)$ underperforms $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$, but outperforms $f^{(\text{BW})}_{1}(t)$. The poor performance further supports our hypothesis that the functional form of $f(t)$ is important when imposing a locality bias.
Table A8: We compare Powerformer’s performance with WCMHA and the Butterworth filter mask $f^{(\text{BW})}_{2}(t)$ on standard time-series datasets for varying decay lengths. The best results are bolded and second best are underlined.
@@
A6.3 Learnable Mask Time Decays
Although we treat the decay time $\alpha$ as a hyperparameter, one may consider learning the optimal decay time by making $\alpha$ learnable. We observe that $\alpha$ monotonically decreases and in some cases flips sign, changing our mask from a power-law decay to a power-law growth. To combat this continuous decrease we applied a learning rate scheduler to limit how far $\alpha$ can change from its initialization. To fairly compare with our hyperparameter results, we initialize $\alpha$ at the same times used in our hyperparameter tuning. We present the aggregate MSE and MAE results in Table A9.
We observe very minor differences between the best hyperparameter results and the best results from learning $\alpha$. During training, we still observe a monotonically decreasing $\alpha$ that slows down only because of the learning rate scheduler. This makes sense since the training algorithm gains access to more information as $\alpha$ decreases and facilitates overfitting. This behavior further supports our claim that $\mathbf{M}^{(\text{C,D})}$ and $\alpha$ act as regularizers.
Table A9: We train Powerformer with $f^{(\text{PL})}(t)$ and learnable time constant and present the results here. The base cases are the best results from Powerformer with constant decay weights: Table A5. The remaining columns designate the initialized time decay.

### v2

@@ Acknowledgements
Figure S17: We present the Powerformer attention score (inset) and weight distributions on the Weather dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents RBCA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,R})}$, and the solid lines represent RBCA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,R})}$.
@@
Figure S18: We present the Powerformer attention score (inset) and weight distributions on the Weather dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents RBCA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,R})}$, and the solid lines represent RBCA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,R})}$.
@@
Figure S19: We present the Powerformer attention score (inset) and weight distributions on the Electricity dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents RBCA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,R})}$, and the solid lines represent RBCA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,R})}$.
@@
Figure S20: We present the Powerformer attention score (inset) and weight distributions on the Electricity dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents RBCA with $f^{(\text{PL})}(t)$ results before applying $\mathbf{M}^{(\text{C,R})}$, and the solid lines represent RBCA with $f^{(\text{PL})}(t)$ results after applying $\mathbf{M}^{(\text{C,R})}$.
S7.2 Similarity Power-Law Mask
@@
Table S6 presents the aggregate MSE and MAE results for the similarity power-law mask ($f^{(\text{SPL})}(t)$) with varying time decays ($\alpha$), forecast lengths, and input sequence lengths. The 512 input sequence generally outperforms the shorter 336 input length. Compared to $f^{(\text{PL})}(t)$, there are more instances in which a 336 input sequence length outperforms. This is likely due to the much faster decay of $f^{(\text{SPL})}(t)$ compared to $f^{(\text{PL})}(t)$. As expected, the best performing $\alpha$ varies between datasets. However, there is much less variation as a function of forecast length and input sequence length when compared to $f^{(\text{PL})}(t)$.
@@
We compare all datasets’ attention score and weight distributions for recency biased attention in Figs S21-S24. Due to the faster decay of $f^{(\text{SPL})}(t)$ we observe stronger bimodal distributions for all datasets than for $f^{(\text{PL})}(t)$, with wider $\mathbf{C}^{(\text{C,R})}$ distributions for ETTm1, ETTm2, and Weather.
@@
Table S6: We compare Powerformer’s performance with the similarity power-law mask $f^{(\text{SPL})}(t)$ on standard time-series datasets for varying decay lengths. The best results are bolded and the second best are underlined.
@@
Figure S21: We present Powerformer’s attention score and weight distributions on the Weather dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents RBCA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,R})}$, and the solid lines represent RBCA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,R})}$.
@@
Figure S22: We present Powerformer’s attention score and weight distributions on the Weather dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents RBCA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,R})}$, and the solid lines represent RBCA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,R})}$.
@@
Figure S23: We present Powerformer’s attention score and weight distributions on the Electricity dataset with a forecast and input length of 96 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents RBCA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,R})}$, and the solid lines represent RBCA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,R})}$.
@@
Figure S24: We present Powerformer’s attention score and weight distributions on the Electricity dataset with a forecast and input length of 720 and 512, respectively. The dotted line represents the reference MHA results, the dashed-dotted line represents RBCA with $f^{(\text{SPL})}(t)$ results before applying $\mathbf{M}^{(\text{C,R})}$, and the solid lines represent RBCA with $f^{(\text{SPL})}(t)$ results after applying $\mathbf{M}^{(\text{C,R})}$.
S7.3 Butterworth Filter: Order 1
@@
Here we evaluate the Butterworth order 1 mask ($f^{(\text{BW})}_{1}(t)$) on the ETT and Weather datasets with a 336 look-back window. The $f^{(\text{BW})}_{1}(t)$ decay is not as steep as $f^{(\text{PL})}(t)$ but has a more gradual decay than $f^{(\text{BW})}_{2}(t)$, as shown in Figs. S4 and 2. This mask offers a middle ground between the step function like $f^{(\text{BW})}_{2}(t)$ and the faster decaying $f^{(\text{PL})}(t)$. Table S7 presents the results for Powerformer with RBCA and $f^{(\text{BW})}_{1}(t)$. We observe that $f^{(\text{BW})}_{1}(t)$ underperforms $f^{(\text{PL})}(t)$, $f^{(\text{SPL})}(t)$ and $f^{(\text{BW})}_{2}(t)$. The poor performance further supports our hypothesis that the functional form of $f(t)$ is important when imposing a locality bias.
@@
Table S7: We compare Powerformer’s performance with RBCA and the order 1 Butterworth Filter mask $f^{(\text{BW})}_{1}(t)$ on standard time-series datasets for varying decay lengths. The best results are bolded and second best are underlined.
@@
S7.4 Butterworth Filter: Order 2
Here we evaluate the Butterworth order 2 mask ($f^{(\text{BW})}_{2}(t)$) on the ETT and Weather datasets with a 336 look-back window. The $f^{(\text{BW})}_{2}(t)$ decay is analogous to the step function, but has a smooth decay to 0 at the end of the window, as shown in Figs. S4. We use $f^{(\text{BW})}_{2}(t)$ as an improved step function to avoid the sharp cutoff. Table S8 presents the results for Powerformer with RBCA and $f^{(\text{BW})}_{2}(t)$. We observe that $f^{(\text{BW})}_{2}(t)$ underperforms $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$, but outperforms $f^{(\text{BW})}_{1}(t)$. The poor performance further supports our hypothesis that the functional form of $f(t)$ is important when imposing a locality bias.
Table S8: We compare Powerformer’s performance with RBCA and the Butterworth filter mask $f^{(\text{BW})}_{2}(t)$ on standard time-series datasets for varying decay lengths. The best results are bolded and second best are underlined.
@@
S8 ABLATION STUDIES
Here, we perform ablation experiments on Powerformer. To separate the effects of the causal mask and the recency bias, we train three Powerformer variants: with MHA, with the causal mask only, and with RBCA. We train and evaluate these variants on the ETT datasets and the Weather dataset for the 96 and 336 prediction lengths. Table S8 shows the corresponding results.
Table S9: Ablation results using vanilla non-causal MHA, causal MHA, and RBCA.
MHA Causal MHA RBCA Metric MSE MAE MSE MAE MSE MAE ETTh1 96 0.370 0.400 0.361 0.390 0.361 0.390 336 0.422 0.440 0.413 0.429 0.406 0.420 ETTh2 96 0.274 0.336 0.270 0.334 0.269 0.334 336 0.331 0.380 0.323 0.380 0.323 0.380 ETTm1 96 0.290 0.342 0.298 0.344 0.285 0.342 336 0.366 0.392 0.370 0.384 0.357 0.385 ETTm2 96 0.165 0.255 0.163 0.253 0.163 0.251 336 0.278 0.329 0.272 0.324 0.273 0.326 Weather 96 0.149 0.198 0.147 0.197 0.147 0.197 336 0.245 0.282 0.243 0.281 0.243 0.279
For both models, we examine the effects of different masks ($f^{(\text{PL})}(t)$, $f^{(\text{SPL})}(t)$, and $f^{(\text{BW})}_{n}(t)$) and various time decays $\alpha$. We do not perform the same amount of experiments for each model and mask pairing due to poor performance. In the remainder of this section, we provide detailed tables of the MSE and MAE scores, as well as plots of the attention scores and weights. In the last subsection (S8.1), we investigate the effects of making $\alpha$ a learnable parameter.
S8.1 Learnable Mask Time Decays
Although we treat the decay time $\alpha$ as a hyperparameter, one may consider learning the optimal decay time by making $\alpha$ learnable. We observe that $\alpha$ monotonically decreases and in some cases flips sign, changing our mask from a power-law decay to a power-law growth. To combat this continuous decrease we applied a learning rate scheduler to limit how far $\alpha$ can change from its initialization. To fairly compare with our hyperparameter results, we initialize $\alpha$ at the same times used in our hyperparameter tuning. We present the aggregate MSE and MAE results in Table S10.
We observe very minor differences between the best hyperparameter results and the best results from learning $\alpha$. During training, we still observe a monotonically decreasing $\alpha$ that slows down only because of the learning rate scheduler. This makes sense since the training algorithm gains access to more information as $\alpha$ decreases and facilitates overfitting. This behavior further supports our claim that $\mathbf{M}^{(\text{C,R})}$ and $\alpha$ act as regularizers.
Table S10: We train Powerformer with $f^{(\text{PL})}(t)$ and learnable time constant and present the results here. The base cases are the best results from Powerformer with constant decay weights: Table S5. The remaining columns designate the initialized time decay.